# GUARDIAN: Safeguarding LLM Multi-Agent Collaborations with Temporal Graph Modeling

### v1

@@ 5.3 Ablation Studies
We investigate two crucial parameters as shown in Figure 6: $\alpha$ controls the balance between structural and attribute reconstruction, while $\gamma$ regulates the compression-relevance trade-off in the information bottleneck (detailed in Appendix A.10). Using GPT-3.5-turbo with 4 agents, optimal performance is achieved when $\alpha\in[0.3,0.5]$ and $\gamma\in[0.001,0.01]$. These ranges suggest that a moderate $\alpha$ effectively balances attribute and structural information, preventing either aspect from dominating reconstruction. The optimal range of $\gamma$ ensures effective compression, preventing excessive information loss while avoiding retention of noise and redundant patterns in multi-agent interactions.
@@ A.5 More Information about Setup
For evaluation metrics, we report model accuracy, anomaly detection rate, and the number of API calls. Model accuracy represents the correctness of final answers produced by the multi-agent system, serving as our primary evaluation metric. The anomaly detection rate measures the system’s ability to identify anomalous patterns during inter-agent interactions [48]. Specifically, we employ a Time-decaying Weighted Anomaly Detection Rate that assigns greater importance to anomalies detected in earlier interaction rounds. This weighting mechanism is implemented through either exponential or linear decay functions, where anomalies in earlier rounds receive higher weights. This approach is particularly valuable for our scenario, as it emphasizes the system’s capability to identify hallucination amplification or anomaly propagation at early stages. For each round, if the detected anomalous node is indeed anomalous (true positive), we assign a score of 1; otherwise, 0. The final anomaly detection rate is calculated as the weighted average across all rounds. API calls serve as a measure of communication efficiency between agents. We use API calls rather than token consumption alone, as token usage shows high variability across different tasks and prompting strategies and thus cannot reliably reflect communication efficiency.
@@ A.8 Additional Experiments on Error Injection and Propagation
Table 8 and  9 present the API call statistics for error injection and propagation. Table A.8 and A.8 show results on the Biographies dataset.
@@ A.10 Details about Ablation Studies
$\gamma$ is derived from the simplification of $\beta$ and $\lambda$. Specifically, starting from the initial total loss function $L_{total}=L_{rec}+\lambda L_{GIB}$, substituting $L_{GIB}$ yields $L_{total}=L_{rec}+\lambda(I(\mathbf{X};\mathbf{Z})-\beta I(\mathbf{Z};\mathbf{Y}))$. Research [45] has shown that the reconstruction loss can be approximated as the negative of mutual information, i.e., $L_{rec}\approx-I(\mathbf{Z};\mathbf{Y})$. After replacing $L_{rec}$ with $L_{res}$, we finally obtain $L_{total}=L_{res}+\frac{\lambda}{1+\beta}I(\mathbf{X};\mathbf{Z})$, where $\gamma=\frac{\lambda}{1+\beta}$.

### v2

@@ 5.3 Ablation Studies
We investigate two crucial parameters as shown in Figure 6: $\alpha$ controls the balance between structural and attribute reconstruction, while $\gamma$ regulates the compression-relevance trade-off in the information bottleneck (More details are presented in Appendix A.10). Using GPT-3.5-turbo with 4 agents, optimal performance is achieved when $\alpha\in[0.3,0.5]$ and $\gamma\in[0.001,0.01]$. These ranges suggest that a moderate $\alpha$ effectively balances attribute and structural information, preventing either aspect from dominating reconstruction. The optimal range of $\gamma$ ensures effective compression, preventing excessive information loss while avoiding retention of noise and redundant patterns in multi-agent interactions.
@@ A.5 More Information about Setup
For evaluation metrics, we report model accuracy, anomaly detection rate, False Discovery Rate (FDR), the number of API calls, and runtime efficiency. Model accuracy represents the correctness of final answers produced by the multi-agent system, serving as our primary evaluation metric. The anomaly detection rate measures the system’s ability to identify anomalous patterns during inter-agent interactions [51]. Specifically, we employ a Time-decaying Weighted Anomaly Detection Rate that assigns greater importance to anomalies detected in earlier interaction rounds. This weighting mechanism is implemented through either exponential or linear decay functions, where anomalies in earlier rounds receive higher weights. This approach is particularly valuable for our scenario, as it emphasizes the system’s capability to identify hallucination amplification or anomaly propagation at early stages. For each round, if the detected anomalous node is indeed anomalous (true positive), we assign a score of 1; otherwise, 0. The final anomaly detection rate is calculated as the weighted average across all rounds. The False Discovery Rate (FDR) directly addresses a critical practical question: “When GUARDIAN flags a node as anomalous and intervenes, what is the probability that this is a false alarm?” Technically, FDR is calculated as the ratio of false positives (nodes incorrectly flagged as anomalous) to the total number of detections, measuring the precision and trustworthiness of our system’s interventions. API calls serve as a measure of communication efficiency between agents. We use API calls rather than token consumption alone, as token usage shows high variability across different tasks and prompting strategies and thus cannot reliably reflect communication efficiency. Runtime efficiency is measured as the average time per question in seconds, evaluating the computational overhead of our approach compared to baseline methods.
@@ A.8 Additional Experiments on Error Injection and Propagation
Table 11 and  12 present the API call statistics for error injection and propagation. Table A.8 and A.8 show results on the Biographies dataset.
@@ A.10 Details about Ablation Studies
The hyperparameter $\gamma$ is derived from the trade-off between reconstruction and information compression. Our total loss function is $L_{\text{total}}=L_{\text{rec}}+\lambda(I(\mathbf{X};\mathbf{Z})-\beta I(\mathbf{Z};\mathbf{Y}))$. Using the approximation $L_{\text{rec}}\approx-I(\mathbf{Z};\mathbf{Y})$ from prior work [48], the objective can be rewritten as:
$L_{\text{total}}\approx\lambda I(\mathbf{X};\mathbf{Z})+(1+\lambda\beta)L_{\text{rec}}$
By normalizing the total loss, we can express the relationship as:
$\frac{L_{\text{total}}}{1+\lambda\beta}\approx\gamma I(\mathbf{X};\mathbf{Z})+L_{\text{rec}},\quad\text{where}\quad\gamma=\frac{\lambda}{1+\lambda\beta}$
Here, $\gamma$ directly controls the weight of the information compression term relative to the reconstruction loss.